# MiniPLM: Knowledge Distillation for Pre-Training Language Models

### v2

@@ Appendix B More Experimental Details
We mostly follow Brown et al. [10] to set the model and learning rate configurations, as summarized in Table 6. The 500M, 1.8B, and 4B teacher models are the officially released Qwen-1.5 checkpoints  [^3] https://huggingface.co/Qwen and the 300M teacher model used in Section 3.5 to analysis the effect of teacher LM sizes Is pre-trained on 200B tokens from our pre-training corpus. We train all the LMs with the AdamW [50] optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a 0.1 weight decay. We set the batch size to 512 and the max sequence length to 1,024, corresponding to 100K total training steps for roughly 50B tokens in Pre-Train w/o KD, SeqKD, and MiniPLM. For MiniLLM and Vanilla KD, we limit the training steps to align their training computation with Pre-Train w/o KD. Specifically, we assume the computation of a forward and backward pass are $2ND$ and $4ND$, respectively, where $N$ is the model size and $D$ is the number of trained tokens. The training steps of MiniLLM and Vanilla KD are listed in Table 7. We linearly warm up the learning rate for 2K steps and apply cosine learning rate decay until 1/10 of the max values. All experiments are conducted on NVIDIA 40G A100 and NVIDIA 32G V100 GPUs.
Model Size $d_{\text{model}}$ $d_{\text{FFN}}$ $n_{\text{layers}}$ $n_{\text{head}}$ $d_{\text{head}}$ learning rate
104M 512 1,408 8 8 64 $6\times 10^{-4}$
200M 768 2,112 12 12 64 $6\times 10^{-4}$
300M 768 2,112 18 12 64 $6\times 10^{-4}$
500M 1,024 2,816 24 16 64 $3\times 10^{-4}$
1.2B 1,536 4,224 24 16 96 $2.5\times 10^{-4}$
@@
Our downstream datasets for evaluation include Hellaswag (HS; 90), LAMBADA (LAM; 61), Winograde (Wino; 43), OpenbookQA (OBQA; 53), ARC-Easy/Challange (ARC-e/c; 16), PIQA [6], SIQA [66], and StoryCloze (Story; 56). We apply the LM-Eval-Harness [24] 444 framework to conduct zero-shot evaluation. We sample 10K documents from the DCLM [45] corpus to construct our test set for language modeling evaluation.
Appendix C More Results
C.1 Computation Scaling Curves of More Sizes
@@
C.2 Test Losses Extrapolation with Scaling Laws
@@
In Table 2, we follow Hoffmann et al. [37] to fit the scaling law curves with the test losses on the DCLM corpus. Then, we used the fitted constants to predict the test losses for 1T and 10T training data in Pre-Train w/o KD and MiniPLM, and that for Vanilla KD when it consumes the same training FLOPs. Specifically, we consider the following power law suggested by Hoffmann et al. [37]:
@@
In Section 3.4, we extrapolate the loss curve of Pre-Train w/o KD with the Data-Constrained Scaling Law [57], which is almost the same as Eq. (15), except that $N_{d}$ represents the total number of tokens in the pre-training corpus repeated 4 times, as suggested by Muennighoff et al. [57]. Therefore, after $N_{d}$ is solved by letting the loss of Pre-Train w/o KD equal that of MiniPLM ($\alpha=0.25$, 4 Eps.), we divide the value by $4$, resulting in a training tokens number of 118B. This means 68B extra training tokens are needed for Pre-Train w/o KD to achieve the performance of MiniPLM using 50B training tokens.
C.3 Difference Sampling with a Proxy Model
@@ Results.
We term this method as $\textsc{MiniPLM}_{\text{prx}}$ and compare its performance with Vanilla KD and MiniPLM in Table 9. We can see that $\textsc{MiniPLM}_{\text{prx}}$ requires much less offline inference computation compared to standard MiniPLM while still maintaining substantial improvement over Vanilla KD.
C.4 Ablation Study
Impact of the Reference Model.
Figure 8: Impact of the reference model size. We use the 1.8B LM as the teacher and the 200M LM as the student. We report the average zero-shot accuracy on the downstream tasks of the LMs trained with MiniPLM and compare it with Vanilla KD.
Figure 9: Impact of the difference sampling ratio $\alpha$. We report the average zero-shot accuracy on the downstream tasks of the LMs trained with MiniPLM, using $\alpha\in[0.3,0.4,0.5,0.6,0.7,0.9]$ and compare it with Vanilla KD.
Difference Sampling Ratio.
C.5 Case Study
We present a case study in Table 10, 11, and 12 to show the instances corresponding to the three parts of data distribution shown in Figure 3(b).

### v3

@@ Appendix B More Experimental Details
We mostly follow Brown et al. (2020) to set the model and learning rate configurations, as summarized in Table 6. The 500M, 1.8B, and 4B teacher models are the officially released Qwen-1.5 checkpoints  [^3] https://huggingface.co/Qwen and the 300M teacher model used in Section 3.5 to analysis the effect of teacher LM sizes Is pre-trained on 200B tokens from our pre-training corpus. We train all the LMs with the AdamW (Loshchilov & Hutter, 2019) optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a 0.1 weight decay. We set the batch size to 512 and the max sequence length to 1,024, corresponding to 100K total training steps for roughly 50B tokens in Pre-Train w/o KD, SeqKD, and MiniPLM. For MiniLLM and Vanilla KD, we limit the training steps to align their training computation with Pre-Train w/o KD. Specifically, we assume the computation of a forward and backward pass are $2ND$ and $4ND$, respectively, where $N$ is the model size and $D$ is the number of trained tokens. The training steps of MiniLLM and Vanilla KD are listed in Table 7. We linearly warm up the learning rate for 2K steps and apply cosine learning rate decay until 1/10 of the max values. All experiments are conducted on NVIDIA 40G A100 and NVIDIA 32G V100 GPUs.
Family Size Vocab. $d_{\text{model}}$ $d_{\text{FFN}}$ $n_{\text{layers}}$ $n_{\text{head}}$ $d_{\text{head}}$ learning rate
Qwen 104M 151,936 512 1,408 8 8 64 $6\times 10^{-4}$
200M 151,936 768 2,112 12 12 64 $6\times 10^{-4}$
300M 151,936 768 2,112 18 12 64 $6\times 10^{-4}$
500M 151,936 1,024 2,816 24 16 64 $3\times 10^{-4}$
1.2B 151,936 1,536 4,224 24 16 96 $2.5\times 10^{-4}$
LLaMA3.1 212M 128,000 768 3,072 12 12 64 $6\times 10^{-4}$
Vocab. $d_{\text{model}}$ $d_{\text{FFN}}$ $n_{\text{layers}}$ conv rank learning rate
Mamba 140M 50,280 768 1,536 24 4 48 $3\times 10^{-3}$
@@
Configurations of KD Across Model Families
The model configurations of LLaMA3.1 (Dubey et al., 2024) and Mamba (Gu & Dao, 2023) for the cross-family distillation experiments are listed in Table 6. The pre-training data and optimization hyper-parameters are the same as the configurations in our main experiments on Qwen, as listed above.
@@
Our downstream datasets for evaluation include Hellaswag (HS; Zellers et al., 2019), LAMBADA (LAM; Paperno et al., 2016), Winograde (Wino; Levesque et al., 2012), OpenbookQA (OBQA; Mihaylov et al., 2018), ARC-Easy/Challange (ARC-e/c; Clark et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), and StoryCloze (Story; Mostafazadeh et al., 2016). We apply the LM-Eval-Harness (Gao et al., 2024) 444 framework to conduct zero-shot evaluation. We sample 10K documents from the DCLM (Li et al., 2024a) corpus to construct our test set for language modeling evaluation.
Appendix C Discussion on Resource Consumption
Time Complexity Space Complexity Actual Time Max Batch Size
Reference LM Training $O(6N_{\text{ref}}D_{\text{ref}})$ $O(3.75N_{\text{ref}})$ 2.6h 32
Teacher LM Inference $O(2N_{\text{tch}}D)$ $O(2N_{\text{tch}})$ 68h 16
Reference LM Inference $O(2N_{\text{ref}}D)$ $O(2N_{\text{ref}})$ 7h 128
Top-k Sampling $O(D)$ 10min -
Table 8: Asymptotic complexity, actual time and memory use of Difference Sampling. $N_{\text{ref}}$ and $N_{\text{tch}}$ are the reference and teacher model sizes. $D$ and $D_{\text{ref}}$ are the sizes of the original corpus and the corpus for training the reference model. “Actual Time” is the running time of each stage. “Max Batch Size” represents the max single GPU batch size, which measures the memory use of each stage.
Time Complexity Space Complexity Actual Time Max Batch Size
Pre-Train $O(6N_{\text{stu}}D)$ $O(3.75N_{\text{stu}})$ 68h 16
Vanilla KD $O(6N_{\text{stu}}D+2N_{\text{tch}}D)$ $O(3.75N_{\text{ref}}+2N_{\text{tch}})$ 139h 8
MiniPLM $O(6N_{\text{stu}}D)$ $O(3.75N_{\text{stu}})$ 68h 16
Table 9: Training-time asymptotic complexity, actual time and memory use of different pre-training methods. $N_{\text{stu}}$ and $N_{\text{tch}}$ are the student and teacher model sizes. $D$ is the size of the pre-training corpus. “Actual Time” is the training time of each method. “Max Batch Size” represents the max single GPU batch size, which measures the memory use of each method. Unlike the setting in Section 3.1, we align the total training steps (token numbers $D$) of the three methods to compare their training-time complexities.
In this section, we discuss the asymptotic complexity, actual runtime, and memory usage of different methods. Table 8 summarizes the offline computation requirements for Difference Sampling, which only need to be performed once for any LMs to be distilled from a specific teacher LM. Table 9 provides the training-time asymptotic complexity, actual runtime, and memory usage for various pre-training methods. For these results, we assume the LMs uses ZeRO-2 (Rajbhandari et al., 2020) and FP16 precision during training to calculate space complexity. The "Actual Time" and "Max Batch Size" metrics are derived from experiments conducted on 8 A100 80GB GPUs. We can see that the inference time of the teacher and reference LMs constitutes the majority of the runtime for Difference Sampling. This cost can be further reduced by leveraging a proxy model for sampling, as detailed in Appendix E. Regarding training-time resource consumption, MiniPLM requires the same time and memory as pre-training without KD, whereas Vanilla KD incurs significantly higher runtime and GPU memory usage due to the online inference of the teacher LM.
Appendix D More Results
@@
D.1 Computation Scaling Curves of More Sizes
D.2 Test Losses Extrapolation with Scaling Laws
@@
In Table 2, we follow Hoffmann et al. (2022) to fit the scaling law curves with the test losses on the DCLM corpus. Then, we used the fitted constants to predict the test losses for 1T and 10T training data in Pre-Train w/o KD and MiniPLM, and that for Vanilla KD when it consumes the same training FLOPs. Specifically, we consider the following power law:
@@
In Section 3.4, we extrapolate the loss curve of Pre-Train w/o KD with the Data-Constrained Scaling Law (Muennighoff et al., 2023), which is almost the same as Eq. (15), except that $N_{d}$ represents the total number of tokens in the pre-training corpus repeated 4 times, as suggested by Muennighoff et al. (2023). Therefore, after $N_{d}$ is solved by letting the loss of Pre-Train w/o KD equal that of MiniPLM ($\alpha=0.25$, 4 Eps.), we divide the value by $4$, resulting in a training tokens number of 118B. This means 68B extra training tokens are needed for Pre-Train w/o KD to achieve the performance of MiniPLM using 50B training tokens.
D.3 Ablation Study
Impact of the Reference Model.
Figure 8: Impact of the reference model size. We use the 1.8B LM as the teacher and the 200M LM as the student. We report the average zero-shot accuracy on the downstream tasks of the LMs trained with MiniPLM and compare it with that of Vanilla KD.
Figure 9: Impact of the difference sampling ratio $\alpha$. We report the average zero-shot accuracy on the downstream tasks of the LMs trained with MiniPLM, using $\alpha\in[0.3,0.4,0.5,0.6,0.7,0.9]$ and compare it with that of Vanilla KD.
Difference Sampling Ratio.
D.4 Correlations between $\log\frac{p(x)}{p_{\text{ref}}(x)}$ and $\log\frac{p(x)}{q_{\theta}(x)}$
$N_{\text{ref}}$ Pearson Spearman Sampling Acc.
45M 0.743 0.774 80.6
104M 0.826 0.840 87.4
200M 0.856 0.879 89.5
Table 11: Correlations between $\log\frac{p(x)}{p_{\text{ref}}(x)}$ and $\log\frac{p(x)}{q_{\theta}(x)}$. $N_{\text{ref}}$ represents the size of the reference model. $q_{\theta}$ is generated from the 500M student LM. We report the Pearson/Spearman Correlation and the sampling accuracy (Sampling Acc.) of using $\log\frac{p(x)}{p_{\text{ref}}(x)}$.
To examine how replacing $q_{\theta}$ with $p_{\text{ref}}$ in Difference Sampling works, we compute the correlations between $\log\frac{p(x)}{p_{\text{ref}}(x)}$ (using 45M, 104M, and 200M LMs to generate $p_{\text{ref}}$) and $\log\frac{p(x)}{q_{\theta}(x)}$ (using the 500M student LM to generate $q_{\theta}$) on a holdout set. Additionally, we calculate the sampling accuracy, which measures how many instances are correctly classified as selected or discarded when using $\log\frac{p(x)}{p_{\text{ref}}(x)}$ and setting the sampling threshold to 0.5, with the selection using $\log\frac{p(x)}{q_{\theta}(x)}$ serving as the ground truth. As shown in Table 11, using $p_{\text{ref}}$ achieves strong correlations with using $q_{\theta}$ and high data sampling accuracy. This supports the rationale for using $p_{\text{ref}}$ in Difference Sampling as described in Section 2.2.
Appendix E Difference Sampling with a Proxy Model
@@ Results.
We term this method as $\textsc{MiniPLM}_{\text{prx}}$ and compare its performance with Vanilla KD and MiniPLM in Table 12. We can see that $\textsc{MiniPLM}_{\text{prx}}$ requires much less inference computation compared to standard MiniPLM while still maintaining substantial improvement over Vanilla KD.
Appendix F Case Study
We present a case study in Table 13, 14, and 15 to show the instances corresponding to the three parts of data distribution shown in Figure 3(b). The explanation of the three parts are list as follows:
  • $p({\bm{x}})\gtrsim p_{\text{ref}}({\bm{x}})$: As shown in Table 13, the instances whose $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}$ value is close to 0 is discarded (Instance #1 and #2). These instances are full of repeated patterns, which is easily learned by both the small and large LMs. Instance #3 is a dialog from a story, where the teacher LM learns slightly well than the student LM, which is retained for the student LM to learn basic language skills. In this way, these easy and common patterns are down-sampled but not fully discarded.
  • $p({\bm{x}})\gg p_{\text{ref}}({\bm{x}})$: As shown in Table 14, the examples where $p({\bm{x}})$ is much larger than $p_{\text{ref}}({\bm{x}})$ is retained. These examples include long and knowledge-intensive documents (Instance #1), in-context learning examples (Instance #2), and high-quality codes (Instance #3). These examples reveals the key large LM’s advantage over the small LM, which should be up-sampled for the student LM to learn.
  • $p({\bm{x}})\lesssim p_{\text{ref}}({\bm{x}})$: As shown in Table 15, the instances where the larger LM learns worse than the small LM are all discarded. We can see that Instance #1 and Instance #2 are irregular noises, which will be harmful to the training of LMs. The small reference LM easily over-fits their surface patterns like symbols and numbers, while a more powerful large LM better recognizes their unnatural features. Instance #3 is a data collection in an academic paper, but lack of contexts, which is useless for LMs to learn the dependency in contexts. Note that these examples only accounts for 6.3% of the pre-training corpus before Difference Sampling, which is a quite small proportion.